OpenStudy (anonymous):

if O is the center of the cirlce ,BC =10 and OC=13 find OA.

OpenStudy (anonymous):

upload the picture you had before please

OpenStudy (anonymous):

ok so from the image since OA goes straight down and BC is perpendicular to OA then we know that OA cuts exactly half the length of BC. so BC length = 10 --> 10/2 = 5 and we know OC = 13 so we can use pythagorean's theorem to find OA. OA^2 + BC^2 = OC^2 so OA^2 + 5^2 = 13^2 OA^2 = 325 OA = 18.0278

OpenStudy (anonymous):

OA^2 = 144

OpenStudy (anonymous):

so what theorem would you use to find what an arc measures?

myininaya (myininaya):

mouse is correct 169-25=144